# Indoor Location Fingerprinting Privacy: A Comprehensive Survey

### v1

@@ 6.5. Overall Comparison
This section discusses the metrics and datasets used in existing ILFPPM. Each ILFPPM considers various aspects of location privacy issues: cryptography-based ILFPPM mostly try to secure the connection channel between the server and the user and mostly lead to heavy computation and communication overheads. However, anonymization-based solutions implement lightweight protocols either on the device or server to obfuscate the localization mechanism. DP-based approaches encompass various solutions, including CDP, LDP, and geo-indistinguishability each addressing different aspects of the privacy norm within DP. Depending on the context, DP can be applied to both the server and the user. FL-based solutions aim to decentralize the learning mechanism for users to preserve privacy. Table 5 provides insights into the application areas and complexity levels of ILFPPM. Each privacy-preserving method, such as cryptographic methods employed primarily during transmission, anonymization implemented on both servers and devices, DP utilized across devices and servers, and FL mainly applied on devices. Additionally, the complexity levels in terms of computational requirements and expertise needed for the deployment of these mechanisms are reported in this Table.
Table 5. Overall comparison on ILFPPM.
ILFPPM Device/Transmission/Server Complexity
Cryptography Transmission High
Anonymization Device/Server Low
Differential Privacy Device/Server Medium
Federated Learning Device Medium
6.5.1. Evaluation metrics
Here, we discuss various metrics used in the different studies to evaluate the ILFPPM. All these metrics and their associated research are mentioned in Tables 1, 2, 3, 4. We categorize all metrics into the three main groups: localization accuracy, Quality of Service (QoS), and privacy metrics.
Localization Accuracy Metrics: The widely used metric in ILFPPM determines the localization accuracy. Location error metrics are crucial for evaluating the accuracy and precision of localization systems in various applications. These metrics are referred to by the following names: Loc Error, Distance Error, RMSE, MSE, MAE, Accuracy, Error Rate, Success Rate, Spatial Loss, RSS Loss, CR area, Area of ASR, floor detection, Kendall’s Tau distance, control value, count, and area. Location Error, Distance Error, Root Mean Square Error (RMSE), Mean Squared Error (MSE), and Mean Absolute Error (MAE) offer a measure of location error. Accuracy, Error Rate, and Success Rate represent the percentage of correctly localized instances, offering a straightforward measure of system performance. Spatial Loss and RSS Loss evaluate the discrepancy between predicted and actual locations within a spatial context, considering the geometric relationships between points. Additionally, metrics such as Cloaked Region (CR) Area, Area of ASR (Area under the Success Rate curve), and Floor Detection accuracy are tailored for specific localization contexts.
Quality of Service (QoS) Metrics: The second group is employed to assess the Quality of Service (QoS) of the ILFPPM. It includes various metrics, such as: QoS Loss, Communication Overhead, Time cost, Run Time, Execution time, Response time, Energy consumption, Entropy, number of messages, Noise Effect, Bandwidth cost, and battery life. QoS loss, measures the degradation in service quality concerning predetermined thresholds, indicating the system’s ability to meet desired performance levels. Computation and communication overhead quantify the additional computational or communication resources required to execute tasks, particularly relevant in cryptography-based approaches, where encryption adds computation and communication overheads without altering the data. Time cost metrics encompass various aspects such as execution time, reflecting the duration of task completion, and run time, which measures the duration of system operation, collectively shaping the system’s responsiveness and real-time performance. The number of messages exchanged during communication sessions provides insights into network traffic and resource utilization, influencing system scalability and efficiency. Noise effect metrics evaluate the impact of signal distortion or interference on data transmission and processing accuracy, essential for robustness in noisy environments.
Privacy Metrics: The last group consists of metrics related to privacy. Privacy metrics are essential for evaluating the effectiveness of privacy-preserving mechanisms in data handling and transmission. These metrics are referred to as Location Privacy, Probabilistic, and Normalized Central Penalty (NCP) in ILFPPM papers. Location privacy metrics evaluate the preservation of location privacy in spatial datasets, considering factors such as spatial granularity and proximity disclosure, crucial for protecting individuals’ location information. Probabilistic metrics for DP-based techniques assess the likelihood of privacy breaches in probabilistic databases, quantifying the probability of inferring sensitive information from query responses while preserving DP guarantees, crucial for ensuring privacy in data analysis and querying scenarios. NCP serves as a quantitative measure of privacy violation, indicating the extent to which sensitive information is disclosed in a dataset, with higher values reflecting increased privacy risks.
6.5.2. Datasets
Given the diversity of proposed methods and the fact that many of them may not be highly applicable to real-world datasets, most papers, even those intended for application to real-world datasets, employ their proposed methods in a simulated testbed. The open-source datasets widely used for indoor localization, even in non-privacy-focused papers, are UJIndoorLoc (Torres-Sospedra et al., 2014) and JUIndoorLoc (Roy et al., 2019). The CSUCY (Konstantinidis et al., 2015), KIOSUCI (Konstantinidis et al., 2015), CRAWDAD (Blinn et al., 2005), PosData (Wang et al., 2018), Geolife, and Gowalla (Hemkumar, 2024) are also public datasets that are used in ILFPPM papers. In some papers, the dataset’s location is not explicitly mentioned; instead, they mention the physical properties of the real-world dataset prepared for the proposed method. Some other datasets are Cyberspace Research Institute (CRI) (Moghtadaiee et al., 2019), Sangmyung University (Kim et al., 2018), Rutgers University (Sadhu et al., 2017), Tampere University (Lohan et al., 2017), Colombia University (Quijano and Akkaya, 2019), University of Helsinki (Nieminen and Järvinen, 2021) and Guangzhou Xinguang shopping mall (Chen et al., 2018) datasets.
7. Discussion & Future Research Directions
@@
Advanced hybrid privacy-preserving techniques: As every privacy-preserving technique has its own advantages and limitations, future research endeavors can be towards the development of advanced hybrid ILFPPM. These hybrid techniques aim to overcome the limitations of individual approaches, providing versatile solutions applicable across various sensor modalities and contextual scenarios. By combining different privacy-preserving strategies, these hybrids offer a comprehensive and adaptable framework for robust protection of user data in diverse indoor environments. In response to the evolving attack landscape, advanced hybrid methods can enhance privacy, defending against more potential attacks from breaching user privacy.
@@
Context-aware privacy and usability: Future research could explore context-aware privacy mechanisms for indoor location data, incorporating user-friendly interfaces. This ensures that users can make informed decisions and manage privacy based on the sensitivity of location data in various contexts. The approach recognizes that privacy preservation is context-dependent and proposes techniques that dynamically adapt to different situations, such as stricter measures in healthcare and more leniency in retail settings. The aim is to develop adaptive privacy solutions capable of addressing concerns across diverse indoor applications.
@@
Scalability: With the proliferation of IoT devices and the exponential growth of data, scalability challenges will need addressing. Advancements in FL can address scalability challenges by improving model architectures, communication efficiency, optimization algorithms, robustness against various attacks, and aggregation methods in large-scale IPS.
Ethical considerations and privacy standards: Future research on indoor location privacy needs to address ethical considerations related to location data use. Simultaneously, efforts should focus on developing privacy standards specifically designed for IPS. This integrated approach aims to ensure responsible and trustworthy deployment of LBS. In the evolving landscape of privacy regulations, such as the GDPR in Europe, new techniques must prioritize compliance, aligning privacy practices with legal requirements to establish a robust legal and ethical framework for indoor location data usage. Researchers should also work on the ethical implications of indoor location data, considering fairness, bias mitigation, and principled service deployment.
@@
Privacy benchmarks and collaboration: Establishing privacy benchmarks and metrics is crucial for fostering collaboration among academia, industry, policymakers, and advocacy groups in the realm of ILFPPM. This collaborative ecosystem creates a shared framework for evaluating and advancing ILFPPM, incorporating research insights, practical implementations, informed policy perspectives, and user-centric advocacy. Privacy benchmarks act as a crucial link, bridging theoretical expertise with real-world applications to stimulate innovation. Policymakers can leverage these benchmarks to align regulations, ensuring a responsive and effective regulatory framework. User-centric principles, ethical considerations, and a commitment to continuous improvement serve as foundational elements in this collaborative effort, emphasizing responsible innovation and user trust as central tenets in the evolution of indoor location technologies.
@@
Multimodal privacy integration: In the complex indoor environment, diverse sensor modalities are frequently employed, each possessing distinct strengths and limitations. Advanced approaches offer the capacity to seamlessly integrate data from these various sensors, culminating in comprehension of the indoor area, all while protecting user privacy. For example, the fusion of Wi-Fi and Bluetooth data with inputs from inertial sensors has the potential to significantly elevate not only the precision of indoor positioning but also the efficacy of privacy preservation mechanisms. This integrated approach empowers IPS to harness the complementary strengths of multiple sensor types, thereby enhancing both accuracy and privacy protection, ultimately affording users a more robust and private experience within indoor spaces.
8. Conclusion

### v2

@@ 5.7.2 Comparative assessment of ILFPPM
Cryptography-based ILFPPM primarily focuses on protecting the privacy of the data being processed, rather than securing the communication channel itself. While secure communication channels, such as those protected by Transport Layer Security (TLS), safeguard against external adversaries, cryptographic protocols in privacy-preserving indoor localization methods operate on top of these secure channels. The primary goal of these cryptographic protocols is to ensure that the data remains private even if the party at the other end of the channel is potentially malicious. However, anonymization-based solutions implement lightweight protocols either on the device or server to obfuscate the localization mechanism. DP-based approaches encompass various solutions, including CDP, LDP, and geo-indistinguishability each addressing different aspects of the privacy norm within DP. Depending on the context, DP can be applied to both the server and the user. FL-based solutions aim to decentralize the learning mechanism for users to preserve privacy. As shown in Table 7, each privacy-preserving method, such as cryptographic methods employed primarily during transmission, anonymization implemented on both servers and devices, DP utilized across devices and servers, and FL mainly applied on devices.
5.7.3 Comprehensive guidelines for employing existing ILFPPM
Comprehensive guidelines for utilizing existing ILFPPM help ensure strong privacy safeguards while maintaining system efficiency:
1) Choosing the right privacy method: If location privacy is a primary concern and the threat comes from semi-trusted adversaries, LDP presents an ideal solution. LDP strikes a strong balance between privacy and performance by ensuring sensitive data remains protected while minimizing computational overhead, making it well-suited for environments where both privacy and efficiency are critical. For situations with extremely sensitive data where protection against fully malicious adversaries is essential, cryptographic techniques such as Homomorphic encryption or Paillier Encryption should be considered. These methods offer robust privacy guarantees by allowing computations on encrypted data without exposing sensitive information. Although resource-intensive, they ensure that data remains secure even in highly adversarial conditions. In cases where privacy must be preserved without transmitting raw location data, FL-based approaches are highly effective. These methods facilitate decentralized data processing, making them advantageous for distributed systems, such as mobile applications or IoT-based indoor localization systems. FL enables collaborative learning across multiple devices without sharing sensitive information, reducing the risk of exposing location data while allowing accurate model training and system functionality.
2) Balancing privacy and performance: When real-time performance is a critical factor, it is advisable to adopt privacy methods that reduce both communication and computational overhead. Techniques such as anonymization and Lightweight LDP tend to be more efficient in these contexts, although they might entail a trade-off by sacrificing some degree of privacy protection. In cases where computational resources and energy consumption are constrained, which is more common in mobile or IoT devices, it is essential to select methods specifically designed for efficiency. Approaches like Federated Knowledge Distillation and AutoEncoders are well-suited for these environments, as they optimize resource usage while still maintaining reasonable levels of privacy. These methods allow for effective data processing and model training without overburdening limited computational capacities, enabling seamless operation in resource-constrained settings.
3) Handling different adversary models: To effectively defend against inference attacks or context-linking attacks, utilizing advanced cryptographic solutions such as Homomorphic encryption or to combine FL with DP are considered. These methods offer robust features that can significantly enhance privacy protection in indoor environments where sensitive data is at risk of exposure. For lightweight systems that need to maintain strong privacy against location-based attacks, techniques such as randomized encoding and random response mechanisms are excellent options. These methods provide sufficient protection while incurring minimal performance costs, making them suitable for applications with limited computational resources.
4) Deploying in real-world systems: FL techniques have demonstrated significant potential for application in real-world environments, such as smart homes, indoor navigation systems, and industrial IoT networks. By distributing computational tasks across multiple devices, these methods effectively reduce the risks associated with data transfer, enhancing both privacy and security. When implementing FL, it is essential to customize the approach to meet the specific requirements of the system in question. For instance, using FedAvg is ideal for scenarios where the data distribution is relatively balanced across participating devices. In contrast, if the data is non-independent and identically distributed (non-IID), which is the case in ILF, utilizing methods like FedProx can help address the challenges posed by this type of data variability. By carefully selecting and adapting FL techniques, we can optimize performance, maintain privacy, and ensure that their systems are robust and efficient in real-world applications.
6 Indoor Localization Applications and Privacy Concerns
6.1 Marketing
Contextually-aware location-based marketing strategies have the potential to significantly boost sales and profits in e-commerce by delivering tailored marketing directly to consumers based on their location, especially in shopping centers, utilizing location-based technology. This approach allows real-time communication between consumers and sellers, enhancing the overall shopping experience. Additionally, it aids businesses in tracking customer behaviors, patterns, and foot traffic, offering personalized services based on mobile devices’ information [9, 8]. However, these strategies bring privacy challenges, as the use of location-based technology for personalized marketing raises concerns about user privacy and potential misuse of real-time location data. Tracking customer behaviors without explicit consent and the extensive collection of personal data through personalized mobile devices also pose ethical and privacy issues. Furthermore, the precision of IPS in localizing mobile devices indoors and outdoors intensifies concerns about intrusive tracking if not appropriately safeguarded.
6.2 Ambient Assisted Living and Disaster Management
Ambient assisted living platforms, relying on precise indoor location tracking through technologies like Bluetooth and various IPS, offer significant benefits for the elderly, ill, or disabled individuals, particularly those with neurodegenerative diseases [128]. These platforms facilitate behavioral tracking, monitoring daily activities, movement patterns, vital signs, and detecting potential dangers such as falls or injuries. Smart houses leverage indoor localization to enhance user experiences, enabling the homeowner to control Wi-Fi network access based on device presence [129]. Beyond homes, location tracking techniques find applications in agriculture for monitoring greenhouses, libraries for book location, parking garages for car monitoring, and warehouses for item tracking [8]. In disaster management, indoor localization can pinpoint the precise location of individuals in peril for efficient rescue efforts [130, 131]. Despite these advantages, privacy concerns arise, including the continuous monitoring of sensitive health-related data, potential invasions of privacy in smart homes, perceived intrusive surveillance in various contexts, and privacy implications in disaster management. Striking a balance between the benefits and privacy protection is crucial, necessitating robust data safeguards and ethical guidelines.
6.3 Health Services and Public Safety
Indoor localization holds immense potential for elevating service standards in healthcare, aiding hospital staff in swiftly locating patients and enabling patients to navigate therapy rooms independently. Doctors can monitor patient mobility, visitors can easily find their patients, and operating rooms can be well-equipped [8, 132]. Smart devices’ positions are traceable, facilitating their quick retrieval. The technology’s future applications include nanosensor-based drug delivery for tumor targeting and hazard identification for disaster mitigation [133]. In law enforcement, Bluetooth beacon-based systems aid quick responses, while in smart buildings, Wi-Fi-enabled alarms guide people to safety. Despite these benefits, concerns arise about continuous tracking compromising patient privacy in healthcare, nanosensor-based drug delivery raising intimate monitoring questions, and surveillance implications in safety management and law enforcement. Striking a balance requires careful ethical and legal considerations in these applications.
During the COVID-19 pandemic, for example, IPSs were employed to monitor and enforce social distancing guidelines in various settings, including hospitals and workplaces. While crucial for public health, these technologies raise significant privacy concerns. The continuous and precise tracking of individuals’ movements within indoor spaces may result in the collection of sensitive location data, revealing interactions, routines, and patterns. This raises potential privacy infringements, and there’s a risk of unauthorized access or misuse, challenging data security and individual autonomy. Addressing this complex challenge requires robust privacy safeguards, transparent data handling practices, and clear regulations to ensure the responsible deployment of IPS for COVID-19 measures while respecting privacy rights and maintaining public trust. The global COVID-19 pandemic has highlighted the critical importance of privacy in indoor localization systems, especially in contexts where accurate location tracking is necessary to enforce social distancing measures. As outlined in [134] advancements in privacy-enhancing technologies were key to ensuring that personal data, particularly location information, should be protected while still allowing for real-time monitoring of individuals in shared spaces.
6.4 Surveillance and Tracking
LBSs extend beyond human tracking, finding significant applications in industrial settings, where the automatic tracking of numerous objects is essential for effective management. This involves real-time control of item locations alongside localization and identification, requiring new Medium Access Control layer protocols to prevent collisions [135]. The prediction of autonomous robot locations, especially in the emerging era of collaborative robots (co-bots), enhances safety and efficiency in various tasks, such as safety and collision avoidance in industrial duties [136, 137, 138]. However, the continuous monitoring of objects and devices in industrial contexts raises privacy concerns for workers, impacting workplace privacy. Anticipating robot locations and monitoring employee interactions could similarly affect privacy. Furthermore, the use of location-based access control and continuous tracking for anti-theft systems raises data security and surveillance issues. Additionally, indoor localization in entertainment and broadcasting may involve tracking event attendees and presenting privacy challenges. Achieving a balance between the advantages of these technologies and privacy protection requires the establishment of robust safeguards and guidelines for responsible use [139].
7 Discussion & Future Research Directions
@@
7.1 Future Research on Cryptographic Techniques for Indoor Location Privacy
Based on what is mentioned in Section 5.3, it is clear that the choice of cryptographic technique as an ILFPPM should be carefully aligned with the specific application needs. For real-time systems or applications with limited computational resources, Paillier encryption offers a good balance between privacy protection and performance, making it a suitable choice for scenarios where the adversary model is semi-trusted and the focus is primarily on location privacy. In contrast, for highly sensitive environments with fully malicious adversaries, where both location and data privacy are critical, the use of Garbled Circuits or multi-layer encryption techniques, despite their computational overhead, provides stronger privacy guarantees. It is also recommended to evaluate these cryptographic methods on real-world datasets to ensure practical applicability. Finally, future research should aim to explore hybrid cryptographic methods that can better balance privacy and efficiency, while also addressing protocol-based attacks that exploit weaknesses in communication frameworks.
7.2 Future Research on Anonymization for Indoor Location Privacy
Privacy-preserving anonymization techniques should provide a balance between efficiency, scalability, and privacy guarantees. Given the evolving threat landscape, it is crucial to integrate multiple anonymization models (e.g., $k$-anonymity, $\ell$-diversity, $t$-closeness) to address both attribute disclosure risks and location privacy. Additionally, applying these techniques to real-world datasets with varying levels of granularity and adversary models would provide more comprehensive insights into their effectiveness. Future research should also explore hybrid anonymization methods that combine randomization, permutation, and spatial techniques to reduce communication overhead without compromising privacy. Finally, attention should be paid to evaluating anonymization methods under more complex attack models, such as semantic and probability-based attacks, to ensure their resilience in diverse operational environments.
7.3 Future Research on DP for Indoor Location Privacy
Future work on DP-based ILFPPM should also aim to refine the balance between privacy protection, data utility, and efficiency. As seen in the evolution from CDP to LDP, decentralized privacy mechanisms show greater potential in protecting sensitive location data in untrusted environments. Researchers should focus on developing hybrid DP methods that combine the strengths of both CDP and LDP to mitigate location and data privacy risks. Additionally, incorporating advanced adversary models, such as semantic and probability-based attacks, into the evaluation of DP methods will provide more comprehensive insights into their resilience. Finally, improving scalability by leveraging edge-cloud collaboration and enhancing run-time efficiency will ensure the practical deployment of these mechanisms in real-world ILF systems.
7.4 Future Research on FL for Indoor Location Privacy
Future work should focus on optimizing FL methods for environments with highly variable data distributions, such as multi-building or city-wide systems. Methods like FedProx and federated knowledge distillation show promise but require further exploration to handle non-IID data effectively in large-scale, heterogeneous systems. While existing works have focused primarily on location privacy and inference attacks, research should explore techniques for defending against emerging attack models like model inversion or MIA. Integrating DP and secure multi-party computation alongside federated techniques could offer more comprehensive defenses. Moreover, given the importance of efficiency metrics like energy consumption and run time, future research should focus on reducing the computational and communication overhead associated with advanced privacy techniques. Developing lightweight FL models that maintain strong privacy guarantees without compromising performance, particularly for edge devices, will be crucial. More studies also need to validate FL-based privacy methods using real-world indoor datasets beyond UJIIndoorLoc and private datasets. A wider variety of scenarios and datasets, such as those from smart homes or healthcare facilities, could better demonstrate the effectiveness and generalizability of the proposed methods.
7.5 Advanced hybrid privacy-preserving techniques
@@
As every privacy-preserving technique has its own advantages and limitations, future research endeavors can be towards the development of advanced hybrid ILFPPM. These hybrid techniques aim to overcome the limitations of individual approaches, providing versatile solutions applicable across various sensor modalities and contextual scenarios. By combining different privacy-preserving strategies, these hybrids offer a comprehensive and adaptable framework for robust protection of user data in diverse indoor environments. In response to the evolving attack landscape, advanced hybrid methods can enhance privacy, defending against more potential attacks from breaching user privacy.
7.6 Context-aware privacy and usability
@@
Future research could explore context-aware privacy mechanisms for indoor location data, incorporating user-friendly interfaces. This ensures that users can make informed decisions and manage privacy based on the sensitivity of location data in various contexts. The approach recognizes that privacy preservation is context-dependent and proposes techniques that dynamically adapt to different situations, such as stricter measures in healthcare and more leniency in retail settings. The aim is to develop adaptive privacy solutions capable of addressing concerns across diverse indoor applications.
7.7 Scalability
@@
With the proliferation of IoT devices and the exponential growth of data, scalability challenges will need addressing. Advancements in FL can address scalability challenges by improving model architectures, communication efficiency, optimization algorithms, robustness against various attacks, and aggregation methods in large-scale IPS.
7.8 Ethical considerations and privacy standards
In response to privacy concerns, many countries have enacted regulations to protect individuals’ data. Notable examples include the General Data Protection Regulation (GDPR) in the European Union [^1] https://gdpr-info.eu/,Health Insurance Portability and Accountability Act (HIPAA) [^2] https://www.hhs.gov/hipaa/for-professionals/privacy/laws-regulations/index.html and the California Consumer Privacy Act (CCPA) [^3] https://oag.ca.gov/privacy/ccpa in the United States, and the Personal Information Protection Law (PIPL) [^4] https://www2.deloitte.com/cn/en/pages/risk/articles/personal-information-protection-law.html in China. These laws aim to give individuals greater control over their data, including location information. Specifically for location privacy, indoor localization models can ensure GDPR compliance by integrating specialized GDPR-based access control systems [140].
Each of these regulations has distinct requirements for handling location data. The GDPR classifies location data as personal data and mandates explicit user consent, data minimization, and the right to erasure. It also enforces privacy-by-design principles for ILF systems. In contrast, the CCPA grants users the right to know, opt out, and request deletion of their location data but does not require explicit opt-in consent. Meanwhile, PIPL classifies location data as sensitive information, requiring justification for collection, explicit consent, and strict cross-border data transfer regulations.
Given the growing legal study surrounding location data privacy, ILF systems must align with these global regulations. Compliance strategies include: (1) implementing anonymization techniques such as k-anonymity, and DP to prevent re-identification [49, 102], (2) restricting access to location data using cryptographic protection and FL [13, 107], (3) minimizing data retention and ensuring compliance with data deletion requests [140], and (4) enhancing transparency and user control by allowing users to manage their data and consent preferences [16]. By integrating these legal standards into ILFPPM, service providers can enhance compliance while mitigating the risks associated with location privacy breaches.
7.9 Privacy benchmarks and collaboration
@@
Establishing privacy benchmarks and metrics is crucial for fostering collaboration among academia, industry, policymakers, and advocacy groups in the realm of ILFPPM. This collaborative ecosystem creates a shared framework for evaluating and advancing ILFPPM, incorporating research insights, practical implementations, informed policy perspectives, and user-centric advocacy. Privacy benchmarks act as a crucial link, bridging theoretical expertise with real-world applications to stimulate innovation. Policymakers can leverage these benchmarks to align regulations, ensuring a responsive and effective regulatory framework. User-centric principles, ethical considerations, and a commitment to continuous improvement serve as foundational elements in this collaborative effort, emphasizing responsible innovation and user trust as central tenets in the evolution of indoor location technologies.
7.10 Multimodal privacy integration
@@
In the complex indoor environment, diverse sensor modalities are frequently employed, each possessing distinct strengths and limitations. Advanced approaches offer the capacity to seamlessly integrate data from these various sensors, culminating in comprehension of the indoor area, all while protecting user privacy. For example, the fusion of Wi-Fi and Bluetooth data with inputs from inertial sensors has the potential to significantly elevate not only the precision of indoor positioning but also the efficacy of privacy preservation mechanisms. This integrated approach empowers IPS to harness the complementary strengths of multiple sensor types, thereby enhancing both accuracy and privacy protection, ultimately affording users a more robust and private experience within indoor spaces.
8 Conclusion